OpenStudy (anonymous):

A negatively charged particle -q is placed at the center of a uniformly charged ring, where the ring has a total positive charge Q as shown in Example 19.5. The particle, confined to move along the x axis, is displaced a small distance x along the axis (x is much smaller than a) and released. Show that the particle oscillates in simple harmonic and calculate its frequency.

OpenStudy (unklerhaukus):

can you find the electric field strength along the x axis?

OpenStudy (anonymous):

 force on particle when it is displaced by a small dist x along the axis is

OpenStudy (anonymous):

\[F=-kqQ/(X ^{2} + R ^{2})^{3/2} \]

OpenStudy (anonymous):

since x is very small therefore neglecting it wrt R and we get \[F=-KqQx/R ^{3} = Ma\]

OpenStudy (anonymous):

since a is propotional to -x therefore it is SHM time period = 2pi/omega where omega = sq root(a/x)

OpenStudy (anonymous):

Our colleague responded but has some observations Note that the electric field produced by a ring uniformly loaded along its axis is given by: \[E=k*x*q/(x^{2}+R^{2})^{3/2} \] R is the ring radius Thus, the electrical force that will act on the charge Q is F = E * Q (along the radial direction as well) \[F=-k*x*q*Q/(x^{2}+R^{2})^{3/2} \] (minus sign because it points in the opposite direction to the displacement) When x << R, we can expand the denominator in series: \[F=-k*x*q*Q/(x^{2}+R^{2})^{3/2} \] \[F -\approx\frac{ k*x*q*Q }{ R^{3} } [(1-\frac{3}{2}\frac{x^{2}}{R^{3}} ]\] Then \[F =-\frac{ k*x*q*Q }{ R^{3} }\] Now this is the resultant force acting on the charge Q, the second law of Newton, \[F=ma\] \[-\frac{k*q*Q}{ R^{3} }x =ma\] In a simple harmonic oscillator (mass m and elastic constant k), the second law of Newton in provides: \[-k*x=m*a\] \[a=\frac{d^{2}x}{d t^{2}}=\omega^{2}x\] with \[\omega=(k/m)^{1/2}\] Comparing equation osclador hamõnico with the equation obtained for the load under the action of the force identified the frequency of oscillation \[\omega=(kqQ/R^{3})^{1/2}\] Hope I helped